OpenStudy (richyw):

find \(\frac{\partial z}{\partial x}\) and \(\frac{\partial z}{\partial y}\) as a functions of \(x, y,\text{ and }z\) assuming that \(z=f(x,y)\) satisfies the equation \[x^{2/3}+y^{2/3}+z^{2/3}=1\]

OpenStudy (anonymous):

Taking a partial derivative. For example: \[\frac{\partial z}{ \partial y}\] means that you derive with respect to y and threat x,z as a constant.

OpenStudy (richyw):

yeah I know that. but I don't really get what this question is asking. there has got to be the chain rule in there.

OpenStudy (klimenkov):

All you need is\[\frac {\partial }{\partial x}x^n=n\cdot x^{n-1}\]try to do this with n=2/3.

OpenStudy (anonymous):

for example if u want to get \(\huge \frac{\partial z}{\partial x} \) u have \(\huge x^{-\frac{1}{3}}+\huge z^{-\frac{1}{3}}\frac{\partial z}{\partial x}=0 \) and now u just need to derive \(\huge z^{-\frac{1}{3}}\) from original equation and put it in the equation above

OpenStudy (klimenkov):

Example: \(x+y+z=1\). Find \(\frac {\partial z}{\partial y}\) when z=f(x,y). You have to derive both parts of equation respect to y:\[\frac {\partial x}{\partial y}+\frac {\partial y}{\partial y}+\frac {\partial z}{\partial y}=\frac {\partial }{\partial y}1\]\[0+1+\frac {\partial z}{\partial y}=0\]\[\frac {\partial z}{\partial y}=-1\]Just make the same thing.